A Bibliography of Publications in American

Mathematical Monthly: 1940–1949

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.05

Title word cross-reference

0 [10, 7]. 1 [10, 7]. 13 [4]. 53 [10]. 8 [10]. Addendum [10]. allied [2]. Amer [10]. conformal [3]. converse [9]. Determinants [10, 7, 1]. dimensional [1]. diophantine [5]. elements [10, 7]. equation [5].

Fermat [9]. Fibonacci [2]. Higher [1]. identities [2]. II [9]. inequality [6]. isoperimetric [6]. J [10]. late [10]. linear [5]. mapping [3]. Math [10]. Monthly [10]. MR [10]. Never [10]. note [5]. numbers [8]. partition [8]. peculiar [4]. primitive [4]. problem [6]. profiles [3]. property [4]. roots [4].

sequence [2]. shaping [3]. squares [8].

theorem [9]. too [10]. trigonometric [2].

use [3].

whose [10, 7]. Williamson [10]. wing [3].

References

Oldenburger:1940:HDD

[1] Rufus Oldenburger. Higher dimensional determinants. American Mathematical Monthly, 47(??):25–33, ???? 1940. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Johnston:1940:FSA

[2] L. S. Johnston. The Fibonacci sequence and allied trigonometric identities. American Mathematical Monthly, 47(??):85– 89, ???? 1940. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Burington:1940:UCM

[3] R. S. Burington. On the use of conformal mapping in shaping wing profiles. American Mathematical Monthly, 57(??):362– 373, ???? 1940. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Beiler:1941:PPP

[4] A. H. Beiler. A peculiar property of the primitive roots of 13. American Mathematical Monthly, 48(??):185–187, ???? 1941.

Lehmer:1941:NLD

[5] D. H. Lehmer. A note on the linear diophantine equation. American Mathematical Monthly, 48(??):240–246, ????

1941. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Adler:1945:IPI

[6] Claire Fisher Adler. An isoperimetric problem with an inequality. American Mathematical Monthly, 52(??):59–69, ???? 1945.

Williamson:1946:DWE

[7] J. Williamson. Determinants whose elements are 0 and 1. American Mathematical Monthly, 53(??):427–434, ???? 1946. See addendum [10].

Lehmer:1948:PNS

[8] D. H. Lehmer. On the partition of numbers into squares. American Mathematical Monthly, 55(??):476–481, ???? 1948. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lehmer:1949:CFT

[9] D. H. Lehmer. On the converse of Fermat’s theorem II. American Mathematical Monthly, 56(??):300–309, ???? 1949. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kaplansky:1995:NTL

[10] Irving Kaplansky. Never too late. addendum to: “Determinants whose elements are 0 and 1” [Amer. Math. Monthly 53 (1946), 427–434, MR 8, 128] by J. Williamson. American Mathematical Monthly, 102(3):259, March 1995. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [7].